OpenStudy (anonymous):

Which of the following best describes the technique used to graph the equation using the slope and y-intercept? y = 5x + 10 A. Mark a point at 5 on the y-axis, go up 10 and left one and mark this point B. Mark a point at 5 on the y-axis, go up 10 and right one and mark this point. C. Mark a point at 10 on the y-axis, go down 5 and right one and mark this point. D. Mark a point at 10 on the y-axis, go up 5 and right one and mark this point

OpenStudy (anonymous):

What does 5 represent and what does 10 represent in the equation? That will take you to the solution.

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Seems so. 10 is the y-intercept, so that is where you start. 5 represents the slope, so as a fraction that is 5/1 (5 up / 1 right). Make sure you are clear on the ideas for the future.

OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation. A. y - 3 = x - 2 B. y - 3 = 5/4 (x - 2) C. y - 8 = 4/5 (x - 6) Could you help me?

OpenStudy (anonymous):

Can you find the slope? Remember the form for pt-slope is (y-y1)=m(x-x1).